# Odd coloring graphs with linear neighborhood complexity

### v1

@@ Abstract
We prove that any class of bipartite graphs with linear neighborhood complexity has bounded odd chromatic number. As a result, if $\mathcal{G}$ is the class of all circle graphs, or if $\mathcal{G}$ is any class with bounded twin-width, bounded merge-width, or a forbidden vertex-minor, then $\mathcal{G}$ is $\chi_{odd}$-bounded.
@@ 1 Introduction
The odd chromatic number $\chi_{odd}(G)$ of a graph $G$ is the minimum number of colors needed to properly color $G$ so that for each non-isolated vertex $v$ of $G$, there exists a color which appears on an odd number of vertices in its open neighborhood $N(v)$. This parameter [^1] Note that in the literature there exists another parameter called odd chromatic number, see e.g. [2]. was recently introduced by Petruševski and Škrekovski [37]. Among the main problems and results concerning this parameter is the conjecture of [37] that all planar graphs have odd chromatic number at most $5$, where $8$ is the best current upper bound [36]. An open problem from [10] states that the odd chromatic number is bounded by the maximum degree plus a constant, where current best results can be found in [11]. From a sparsity point of view, classes of bounded expansion have bounded odd chromatic number. Indeed, they even have bounded proper conflict free-chromatic number [27, 33] and bounded strong odd chromatic number [38]; see the end of this section for the relevant definitions.
In the current paper, we prove that any bipartite graph of linear neighborhood complexity has bounded odd chromatic number. There are many examples of graphs with linear neighborhood complexity. In particular, classes of graphs with good model-theoretic properties often have linear neighborhood complexity; see, for example, [3, 4, 20, 40]. The neighborhood complexity of a graph $G$, denoted by $\eta_{G}(m)$, is the maximum over all $m$-element subsets $A\subseteq V(G)$ of |$|\{N(v)\cap A:v\in V(G)\}|$. Thus, the neighborhood complexity counts the number of different neighborhoods that can occur within an $m$-vertex set.
@@
For any integer $r$, there exists an integer $k=k(r)$ so that if $G$ is a bipartite graph with neighborhood complexity $\eta_{G}(m)\leqslant r\cdot m$ for every positive $m$, then $\chi_{odd}(G)\leqslant k$.
The main ingredient in our proof of Theorem 1.1 is Haussler’s Shallow Packing Lemma [26].
We obtain the following corollary of Theorem 1.1 using a technique from [29] to reduce to the case of bipartite graphs. A graph class is $\chi_{odd}$-bounded if there exists a function $f$ so that any graph in the class with clique number $\omega$ has odd chromatic number at most $f(\omega)$.
@@
If $\mathcal{G}$ is any class of graphs with bounded twin-width, bounded merge-width, or a forbidden vertex-minor, then $\mathcal{G}$ is $\chi_{odd}$-bounded.
Our proof of Corollary 1.2 uses the fact that classes with bounded twin-width [5, 9, 39], bounded merge-width [3], or a forbidden vertex-minor [13] are $\chi$-bounded. (The chromatic number of a graph $G$ is denoted by $\chi(G)$, and a class of graphs is $\chi$-bounded if there exists a function $f$ so that any graph $G$ in the class with clique number $\omega$ has $\chi(G)\leqslant f(\omega)$.)
Next, we use a connection between the Matroid Growth Rate Theorem [22] and neighborhood complexity in order to obtain the following corollary of Theorem 1.1. This connection was recently observed by the authors [14], who used it in order to find the unavoidable cosimple $\textsf{GF}(q)$-representable matroids of large girth.
@@
An interesting common special case of Corollary 1.2 and Corollary 1.3 is that of (bipartite) circle graphs. A graph is a circle graph if there exists a collection of chords on a circle so that the vertices are in bijection with the chords, and two vertices are adjacent if and only if their chords have non-empty intersection. Bouchet [8] characterized circle graphs by their forbidden vertex-minors. Additionally, circle graphs with bounded clique number have bounded twin-width [28]. On the other hand, bipartite circle graphs are the fundamental graphs of planar graphs [16]. We also note that there are several direct proofs that circle graphs are $\chi$-bounded; see [12, 15, 25, 30].
Hence, the above results settle [29, Question 3.9], which asked whether circle graphs are $\chi_{odd}$-bounded. Using a technique completely independent of the previous results, we further improve the bound on the odd chromatic number of bipartite circle graphs to 98. This is a more explicit bound, even if it is far from the lower bound; see 5.3.
@@
We obtain this result as a corollary of the more general Theorem 5.2, which can be seen as a “one-sided” odd-coloring result for the fundamental graph of a graphic matroid. We believe that Theorem 5.2 is interesting on its own and propose a possible strengthening; see 5.4.
[…]
@@
In Section 2, we introduce Haussler’s Shallow Packing Lemma and some preliminaries about matroids. We prove Theorem 1.1 in Section 3. In Section 4, we prove the two corollaries of Theorem 1.1 (Corollary 1.2 and Corollary 1.3). Finally, in Section 5, we use a different technique to find a concrete bound on the odd chromatic number of bipartite circle graphs and propose a couple of open problems.
@@ 2 Preliminaries
We write $K_{t}$ for the complete graph with $t$ vertices, $M(G)$ for the cycle matroid of a graph $G$, and $M^{*}$ for the dual of a matroid $M$. We write $U_{2,q}$ for the $q$-element line, and, more generally, $U_{t,q}$ for the uniform matroid with $q$ elements and rank $t$. That is, $U_{t,q}$ is the $q$-element matroid where the circuits are the sets of size $t+1$. Given a graph $G$ and a vertex $v$, we write $N_{G}(v)$ (or just $N(v)$ when the graph is clear from the context) for the open neighborhood of $v$. That is, we do not consider $v$ to be in its neighborhood $N(v)$.
@@
The following version of Haussler’s Shallow Packing Lemma from [26] is stated as [19, Lemma 2.2], for instance. We just require the case that $d=1$.
@@
Let $M$ be a matroid. Given a basis $B$ of $M$, the fundamental circuit of an element $e\in E(M)\setminus B$ with respect to $B$ is the unique circuit of $M$ which is contained in $B\cup\{e\}$. The fundamental graph $\mathcal{F}(M,B)$ is the bipartite graph with sides $B$ and $E(M)\setminus B$ where each element $e\in E(M)\setminus B$ is adjacent to the other elements in its fundamental circuit. Given a graph $G$, the bases of the cycle matroid of $G$ are the edge sets of spanning forests; a spanning forest is a maximal acyclic subgraph of $G$. If $T$ is a spanning forest of $G$, then the fundamental graph of $G$ with respect to $T$ is the graph $\mathcal{F}(M(G),E(T))$. For convenience, we denote this graph by $\mathcal{F}(G,T)$.
Bouchet [7] introduced fundamental graphs for binary matroids and showed how they behave with respect to taking matroid minors. However, since Bouchet used different (more general) terminology, we refer the reader to Oum [35] for a nice introduction to fundamental graphs of binary matroids. For general matroids, see for instance [24].
Lemma 2.2 ([24, Proposition 3.1(ii)]).
@@
For a matroid $M$ and element $e\in E(M)$, we write $M\setminus e$ for the matroid obtained from $M$ by deleting $e$ and $M/e$ for the matroid obtained from $M$ by contracting $e$. For a graph $G$ and a vertex $v\in V(G)$, we write $G-v$ for the graph obtained from $G$ by deleting $v$. Consider deleting a vertex from the non-basis side of a fundamental graph; this corresponds to just deleting that element from the matroid. Thus, we obtain the following lemma, using Lemma 2.2 and the fact that for any matroid $M$ and element $e$, we have $M^{*}\setminus e=(M/e)^{*}$.
@@
In Section 4, we use the Growth Rate Theorem of Geelen and Whittle [22] to prove that the fundamental graphs of certain matroids have linear neighborhood complexity (see Lemma 4.1). The Growth Rate Theorem of Geelen and Whittle [22] says the following. (We remark that there is also a more general Growth Rate Theorem of Geelen, Kung, and Whittle [21].)
@@
Note that graphic matroids exclude the line $U_{2,4}$ as a minor. We remark that Nelson, Norin, and Omana [34] have recently improved the bound on $g(t,q)$ in Theorem 2.4 to a single exponential function.

### v2

@@ Abstract
We prove that any class of graphs with linear neighborhood complexity has bounded improper odd chromatic number. As a result, if $\mathcal{G}$ is the class of all circle graphs, or if $\mathcal{G}$ is any class with bounded twin-width, bounded merge-width, or a forbidden vertex-minor, then $\mathcal{G}$ is $\chi_{\mathrm{o}}$-bounded.
@@ 1 Introduction
The odd chromatic number $\chi_{\mathrm{o}}(G)$ of a graph $G$ is the minimum number of colors needed to properly color $G$ so that for each non-isolated vertex $v$ of $G$, there exists a color which appears on an odd number of vertices in its open neighborhood $N(v)$. This parameter [^1] Note that in the literature there exists another parameter called odd chromatic number, see e.g. [1]. was introduced by Petruševski and Škrekovski [51]. Among the main problems and results concerning this parameter is the conjecture of [51] that all planar graphs have odd chromatic number at most $5$, where $8$ is the best current upper bound [50]. Related classes that have been studied are $1$-planar graphs [44], IC-planar graphs [49], toroidal graphs [57], graphs of bounded thickness [38], and graphs of bounded maximum average degree [12, 14, 58]. Another interesting conjecture in the area, due to Caro, Petruševski and Škrekovski [9], states that the odd chromatic number of a graph is at most its maximum degree plus a constant. The current best results can be found in [15].
While Cranston [14] constructed graphs with maximum average degree at most 4 and unbounded odd chromatic number, there are other more restrictive notions of sparsity that do bound the odd chromatic number. In particular, classes of bounded expansion have bounded odd chromatic number; indeed, they even have bounded proper conflict free-chromatic number [35, 43] and bounded strong odd chromatic number [52]. See the end of this section for the relevant definitions.
A convenient relaxed notion is the improper odd chromatic number $\chi_{\mathrm{io}}(G)$ of a graph $G$. It is the minimum number of colors needed to (possibly improperly) color the vertices of $G$ so that for each non-isolated vertex $v$ of $G$, there exists a color which appears on an odd number of vertices in its open neighborhood $N(v)$. Since a product coloring of an improper odd coloring and a proper coloring yields a proper odd coloring, we obtain the following. We denote the chromatic number of a graph $G$ by $\chi(G)$.
Remark 1.1.
For every graph $G$ we have $\chi_{\mathrm{o}}(G)\leqslant\chi(G)\chi_{\mathrm{io}}(G)$.
In the current paper, we prove that any graph of linear neighborhood complexity has bounded improper odd chromatic number, which then yields bounds on the odd chromatic number once the chromatic number is bounded. There are many examples of graphs with linear neighborhood complexity. In particular, classes of graphs with good model-theoretic properties often have linear neighborhood complexity; see, for example, [2, 3, 28, 54]. The neighborhood complexity of a graph $G$, denoted by $\eta_{G}(m)$, is the maximum over all $m$-element subsets $A\subseteq V(G)$ of $|\{N(v)\cap A:v\in V(G)\}|$. Thus, the neighborhood complexity counts the number of different neighborhoods that can occur within an $m$-vertex set.
@@
For any integer $r$, there exists an integer $k=k(r)$ so that if $G$ is a graph with neighborhood complexity $\eta_{G}(m)\leqslant r\cdot m$ for every positive $m$, then $G$ can be (improperly) $k$-colored such that for each non-isolated vertex $v$, there exists a color which appears an odd number of times in its neighborhood $N(v)$.
The main ingredient in our proof of Theorem˜1.2 is Haussler’s Shallow Packing Lemma [34].
We obtain the following corollary of Theorem˜1.2 using Remark˜1.1 and, in the case of vertex-minors, a technique from [37]. A graph class is $\chi_{\mathrm{o}}$-bounded [^2] Note that in [37] this was defined via the existence of $f$ such that $\chi_{\mathrm{o}}(G)\leqslant f(\chi(G))$ for all $G$ in the class. Since the classes we consider are $\chi$-bounded, both definitions coincide for the results in the present paper. if there exists a function $f$ so that any graph in the class with clique number $\omega$ has odd chromatic number at most $f(\omega)$.
@@
If $\mathcal{G}$ is any class of graphs with bounded merge-width or with a forbidden vertex-minor, then $\mathcal{G}$ is $\chi_{\mathrm{o}}$-bounded.
We note that every class of bounded twin-width also has bounded merge-width [22, Theorem 1.4]. So Corollary 1.3 implies that classes of bounded twin-width are $\chi_{\mathrm{o}}$-bounded. The chromatic number of graphs with bounded twin-width has been well-studied; see [4, 8, 53]. Also see [5] for the original motivation of the parameter.
Our proof of Corollary˜1.3 uses the fact that classes with bounded merge-width [2] or a forbidden vertex-minor [17] are $\chi$-bounded. (A class of graphs is $\chi$-bounded if there exists a function $f$ so that any graph $G$ in the class with clique number $\omega$ has $\chi(G)\leqslant f(\omega)$.)
We also use a connection between the Matroid Growth Rate Theorem [30] and neighborhood complexity in order to obtain the following corollary of Theorem˜1.2 for a class of bipartite graphs. This connection was recently observed by the authors [18], who used it in order to find the unavoidable cosimple $\textsf{GF}(q)$-representable matroids of large girth.
@@
An interesting common special case of Corollary˜1.3 and Corollary˜1.4 is that of bipartite circle graphs. A graph is a circle graph if there exists a collection of chords on a circle so that the vertices are in bijection with the chords, and two vertices are adjacent if and only if their chords have non-empty intersection. Bouchet [7] characterized circle graphs by their forbidden vertex-minors. Additionally, circle graphs with bounded clique number have bounded merge-width [22, 36]. We also note that there are several direct proofs that circle graphs are $\chi$-bounded; see [16, 19, 33, 39]. Moreover, bipartite circle graphs are the fundamental graphs of planar graphs [27]. Using a technique from [37], we apply our results about bipartite circle graphs to prove that all circle graphs are $\chi_{\mathrm{o}}$-bounded. This settles [37, Question 3.9], which asked whether circle graphs are $\chi_{\mathrm{o}}$-bounded.
Based on the fact that bipartite circle graphs are the fundamental graphs of planar graphs, we also develop another strategy to prove $\chi_{\mathrm{o}}$-boundedness. In the case of circle graphs, this strategy gives a much better bound.
For this result, recall that the star-arboricity of a graph $G$ is the smallest integer $\ell$ such that $E(G)$ can be decomposed into $\ell$ star forests (that is, forests where each component has diameter at most $2$). By the Kostochka–Thomason bound [40, 41, 55, 56], for any integer $t$, every graph with no $K_{t}$ minor has star-arboricity $O(t\sqrt{\log t})$. For each integer $t$, we write $\ell(t)$ for the largest star-arboricity of any graph with no $K_{t}$-minor. Given a graph $G$, a spanning forest $T$ of $G$, and an edge $f\in E(T)$, note that deleting $f$ from $T$ yields exactly two components. The fundamental cut of $T$ with respect to $f$ is the set of all edges in $E(G)$ which join these two components. We can then prove the following “one-sided” coloring result.
Theorem 1.5.
For any integer $t$, there exists an integer $k\leqslant 16\ell(t-1)+1$ so that if $G$ is a graph with no $K_{t}$ minor and $T$ is a spanning forest of $G$, then there exists a coloring $\phi:E(G)\setminus E(T)\rightarrow\{1,\ldots,k\}$ so that for each non-bridge $f\in E(T)$, there exists a color $i\in\{1,\ldots,k\}$ which occurs an odd number of times on the fundamental cut of $T$ with respect to $f$.
Theorem˜1.5 can be seen as a one-sided coloring result in the fundamental graph of the graphic matroid of $G$, which is a matroid $M$ without $U_{2,4}$, $F_{7}$, $F^{*}_{7}$, $M^{*}(K_{5})$, $M^{*}(K_{3,3})$, or $M(K_{t})$ as a minor, where $F_{7}$ is the Fano matroid and $K_{3,3}$ is the complete bipartite graph with sides of size $3$. We believe that Theorem 1.5 is interesting on its own and propose a possible strengthening; see ˜5.3. As an application of Theorem˜1.5, we further improve the bound on the odd chromatic number of bipartite circle graphs to 98. This is a more explicit bound, even if it is far from the lower bound; see ˜5.2.
@@
Recall that our proof that circle graphs are $\chi_{\mathrm{o}}$-bounded reduces to the case of bipartite circle graphs using the technique from [37]. We note that this type of approach is necessary in the following sense.
Proposition 1.7.
There are circle graphs of arbitrarily high improper odd chromatic number.
Related results and parameters
We now give a brief survey of how various versions of the chromatic number are related.
Proper coloring parameters.
The odd chromatic number was introduced as a relaxation of the proper conflict-free chromatic number $\chi_{\mathrm{pcf}}(G)$. Here for each non-isolated vertex $v$ of $G$, there must exist a color which appears on exactly one vertex in its open neighborhood $N(v)$. This parameter was introduced by Fabrici, Lužar, Rindošová, and Soták [24]. Both parameters can be strengthened by requiring that every color in the open neighborhood of $v$ appears an odd number of times or exactly once, respectively. This leads to the strong odd chromatic number $\chi_{\mathrm{so}}(G)$ introduced in [42], and the squared chromatic numer, i.e. the chromatic number $\chi(G^{2})$ of $G^{2}$ the square of $G$, where different vertices of $G^{2}$ are connected if their distance in $G$ is at most $2$, see [13] for a recent account. If $\Delta(G)$ denotes the maximum degree of $G$, then from the above definitions, we immediately get the following inequalities.
$\omega(G)\leqslant\chi(G)\leqslant\chi_{\mathrm{o}}(G)\leqslant\chi_{\mathrm{so}}(G),\chi_{\mathrm{pcf}}(G)\leqslant\chi(G^{2})\leqslant\Delta(G)^{2}+1.$ (1)
The rightmost inequality in (1) can also be reversed in some sense, since we trivially have $\chi(G^{2})\geqslant\Delta(G)+1$ for all graphs $G$. However, all remaining inequalities in (1) can be arbitrarily far apart: Petruševski and Škrekovski [51] show that there is no function $f$ such that $\chi_{\mathrm{o}}(G)\leqslant f(\chi(G))$ for all $G$. It is also known [37] that there is no function $f$ such that $\chi_{\mathrm{pcf}}(G)\leqslant f(\chi_{\mathrm{o}}(G))$ for all $G$. Further, since $\Delta(G)<\chi(G^{2})$, any graph class with unbounded degree but bounded $\chi_{\mathrm{pcf}}$ shows that there is no function $f$ such that $\chi(G^{2})\leqslant f(\chi_{\mathrm{pcf}}(G))$ for all $G$, e.g. complete bipartite graphs. Kwon and Park [42] note that there is no function $f$ such that $\chi(G^{2})\leqslant f(\chi_{\mathrm{so}}(G))$ for all $G$. Also, there is no function $f$ such that $\chi_{\mathrm{so}}(G)\leqslant f(\chi_{\mathrm{o}}(G))$ for all graphs $G$, see [31]. Just for completeness, let $G$ be obtained from $K_{n}$ by subdividing each edge once and then attaching one vertex incident to all others. Then $\chi_{\mathrm{pcf}}(G)\leqslant 4$ but $\chi_{\mathrm{so}}(G)\geqslant n$. On the other hand, if $G$ is the bipartite graph with bipartition $(X,Y)$ where ${|X|-1}\choose 2$} is odd and for every $A\subseteq X$ with $|A|=3$ there is exactly one $v\in Y$ with $N(v)=A$, then $\chi_{\mathrm{so}}(G)=2$ but $\chi_{\mathrm{pcf}}(G)\geqslant\frac{|X|}{2}$. So $\chi_{\mathrm{so}}$ and $\chi_{\mathrm{pcf}}$ are not tied together by any function.
Hence, one can wonder if for a class of graphs there is a function that upper bounds one of the above coloring parameters in terms of the clique number, yielding various notions of boundedness. As mentioned in the beginning of the introduction, classes of bounded expansion are $\chi_{\mathrm{pcf}}$-bounded [35, 43] as well as $\chi_{\mathrm{so}}$-bounded [52].
To see the relevance of our results, note that circle graphs are neither $\chi_{\mathrm{pcf}}$-bounded [37] nor $\chi_{\mathrm{so}}$-bounded [31]. Hence they highlight the difference of $\chi_{\mathrm{o}}$ when compared to $\chi_{\mathrm{pcf}}$ and $\chi_{\mathrm{so}}$. Further, we can see that also the classes we consider behave different than other closely related classes. Namely, while the class of permutation graphs, a subclass of circle graphs, is $\chi_{\mathrm{pcf}}$-bounded, several other superclasses of permutation graphs such as convex bipartite graphs, grid-intersection graph, and comparability graphs of 3-dimensional posets are not $\chi_{\mathrm{o}}$-bounded, see [37].
Improper colorings.
Each of the above coloring parameters can be relaxed by dropping the requirement of the coloring being proper. The proper conflict-free colorings were actually motivated by (improper) conflict-free chromatic number $\chi_{\mathrm{cf}}$ introduced by Pach and Tardos [48]. This came naturally from conflict-free colorings of hypergraphs studied by Even, Lotker, Ron, and Smorodinsky [23] in relation to the frequency assignment problem in cellular networks. The latter were later relaxed by Cheilaris, Keszegh, and Pálvölgyi to odd colorings for hypergraphs in [10]. From this point of view the improper odd chromatic number $\chi_{\mathrm{io}}(G)$ arises naturally and has been considered in [11, 60]. The improper strong odd chromatic number $\chi_{\mathrm{iso}}(G)$ was introduced in [31]. The improper version of the squared chromatic number is the chromatic number $\chi(G^{\sharp 2})$ of the exact square $G^{\sharp 2}$ of $G$, where vertices of $G^{\sharp 2}$ are connected if their distance in $G$ is $2$, see [25]. The improper variant of a parameter is always a lower bound for the proper variant. As mentioned in Remark˜1.1 a product coloring yields that $\chi_{\mathrm{o}}(G)\leqslant\chi_{\mathrm{io}}(G)\chi(G)$. Similarly one gets $\chi_{\mathrm{pcf}}(G)\leqslant\chi_{\mathrm{cf}}(G)\chi(G)$ and $\chi(G^{2})\leqslant\chi(G^{\sharp 2})\chi(G)$. However, there exists no function $f$ such that that $\chi_{\mathrm{so}}(G)\leqslant f(\chi_{\mathrm{iso}}(G),\chi(G))$ for all $G$, see [31].
@@
In Section˜2, we introduce Haussler’s Shallow Packing Lemma and some preliminaries about matroids. We prove Theorem˜1.2 in Section˜3. In Section˜4, we prove the two corollaries of Theorem˜1.2 (Corollary˜1.3 and Corollary˜1.4). Finally, in Section˜5, we use a different technique (Theorem˜1.5) to find a concrete bound on the odd chromatic number of bipartite circle graphs (Corollary˜1.6). We also prove Proposition 1.7 and propose a couple of open problems.
@@ 2 Preliminaries
We introduce some basic notions of matroids and graphs, see [47, 59] and [20] for books on these topics.
A matroid is a pair $M=(E,\mathcal{B})$ of a finite ground set $E$ and a collection of bases $\mathcal{B}$. These are subsets of $E$ such that for all $B,B^{\prime}\in\mathcal{B}$ and $e\in B\setminus B^{\prime}$ there exists a $e^{\prime}\in B^{\prime}\setminus B$ such that $B\setminus\{e\}\cup\{e^{\prime}\}\in\mathcal{B}$. The dual matroid of $M$ is $M^{*}=(E,\{E\setminus B\mid B\in\mathcal{B}\})$. A subset $C\subseteq E$ is a circuit of $M$ is it is contained in no basis of $M$, and it is a cocircuit of $M$ is it is a circuit of $M^{*}$. The girth and cogirth of a matroid are the size of its smallest circuit and cocircuit, respectively. If the girth and corgirth is larger than two, we say that $M$ is simple and cosimple, respectively. For an element $e\in E$, we write $M\setminus e$ for the matroid obtained from $M$ by deleting $e$, i.e., its bases are the largest independent sets of $M$ in $E\setminus\{e\}$. Furthermore, denote by $M/e$ the matroid obtained from $M$ by contracting $e$, i.e, $(M^{*}\setminus e)^{*}$. If $M^{\prime}$ can be obtained from $M$ by a sequence of contractions and deletions, we say that $M^{\prime}$ is a minor of $M$. We write $U_{2,q}$ for the $q$-element line, and, more generally, $U_{t,q}$ for the uniform matroid with $q$ elements and rank $t$. That is, $U_{t,q}$ is the $q$-element matroid where the circuits are the sets of size $t+1$. We say that $M$ is $\textsf{GF}(q)$-representable if there exists a $\textsf{GF}(q)$-vector space whose bases correspond to $M$. IF $q=2$, then $M$ is binary.
Given a graph $G$, the bases of the graphic matroid $M(G)$ of $G$ are the edge sets of spanning forests; a spanning forest is a maximal acyclic subgraph of $G$. Given a graph $G$ and a vertex $v$, we write $N_{G}(v)$ (or just $N(v)$ when the graph is clear from the context) for the open neighborhood of $v$. That is, we do not consider $v$ to be in its neighborhood $N(v)$. For a graph $G$ and a vertex $v\in V(G)$, we write GvG-v for the graph obtained from $G$ by deleting $v$. We write $K_{t}$ for the complete graph with $t$ vertices.
@@
The following version of Haussler’s Shallow Packing Lemma from [34] is stated as [26, Lemma 2.2], for instance. The original statement contains a parameter $d$, which for our purposes can be substituted by $1$, yielding the following.
@@
Let $M$ be a matroid. Given a basis $B$ of $M$, the fundamental circuit of an element $e\in E(M)\setminus B$ with respect to $B$ is the unique circuit of $M$ which is contained in $B\cup\{e\}$. The fundamental graph $\mathcal{F}(M,B)$ is the bipartite graph with sides $B$ and $E(M)\setminus B$ where each element $e\in E(M)\setminus B$ is adjacent to the other elements in its fundamental circuit. If $T$ is a spanning forest of $G$, then the fundamental graph of $G$ with respect to $T$ is the graph $\mathcal{F}(M(G),E(T))$. For convenience, we denote this graph by $\mathcal{F}(G,T)$.
Bouchet [6] introduced fundamental graphs for binary matroids and showed how they behave with respect to taking matroid minors. However, since Bouchet used different (more general) terminology, we refer the reader to Oum [46] for a nice introduction to fundamental graphs of binary matroids. For general matroids, see for instance [32]. The following is well-known and easy to see.
Lemma 2.2.
@@
Consider deleting a vertex from the non-basis side of a fundamental graph; this corresponds to just deleting that element from the matroid. Thus, we obtain the following lemma, using Lemma˜2.2 and the fact that for any matroid $M$ and element $e$, we have $M^{*}\setminus e=(M/e)^{*}$.
@@
In Section˜4, we use the Growth Rate Theorem of Geelen and Whittle [30] to prove that the fundamental graphs of certain matroids have linear neighborhood complexity (see Lemma˜4.1). The Growth Rate Theorem of Geelen and Whittle [30] says the following. (We remark that there is also a more general Growth Rate Theorem of Geelen, Kung, and Whittle [29].)
@@
Note that graphic matroids exclude the line $U_{2,4}$ as a minor. We remark that Nelson, Norin, and Omana [45] have recently improved the bound on $g(t,q)$ in Theorem˜2.4 to a single exponential function.